OpenStudy (anonymous):

Solve. 4x + 2 ≥ –6 A. x ≤ –2 B. x ≤ –1 C. x ≥ –2 D. x ≥ –1

OpenStudy (anonymous):

that is correct but now which inequality sign do you use

OpenStudy (anonymous):

the answer is c

OpenStudy (anonymous):

not really ^

OpenStudy (anonymous):

it's negative one

OpenStudy (anonymous):

I'm trying to figure out which sign to use...

OpenStudy (anonymous):

ok take your time, do you want help?

OpenStudy (acxbox22):

it is c

OpenStudy (anonymous):

\[4x-2=-6 4x=-6+2 x=-4/4 x=-1 \]

OpenStudy (anonymous):

no acxbox

OpenStudy (anonymous):

yea definately not c

OpenStudy (anonymous):

it's D

OpenStudy (anonymous):

C is the correct answer !

OpenStudy (anonymous):

WTF!!!???

OpenStudy (anonymous):

lol why does everyone think its c?

OpenStudy (anonymous):

because it's the right answer duhh !

OpenStudy (acxbox22):